# Yong Liu    Guo Qin    Zhiyuan Shi    Zhi Chen    Caiyin Yang    Xiangdong Huang    Jianmin Wang    Mingsheng Long

### v1

@@ 4.1.3 TimeFlow Loss
Given representations $\{\mathbf{h}_{i}\}$ extracted by the last layer of the Transformer, we aim to generate length-$F$ predictions $\widehat{\mathbf{y}}_{i}=\widehat{x}_{1+iP,F+iP}$ at each position $i$ via our autoregressive model. While a small input patch size can accommodate more high-frequency data, a large output patch size can empirically lead to better results (Das et al., 2023b). Therefore, we adopt $F>P$ for multi-patch predictions in our models.
@@
where $\mathbf{y}_{i}\in\mathbb{R}^{F}$ is the groundtruth value and $\mathbf{y}_{i}^{(0)}$ is a $d$-dimensional Gaussian noise, $t$ is sampled from $\mathcal{U}[0,1]$, and $\mathbf{y}_{i}^{(t)}=t\mathbf{y}_{i}+(1-t)\mathbf{y}_{i}^{(0)}$ constructed by the conditional optimal-transport path. It is important to note that the conditional representation $\mathbf{h}_{i}$ differs from the conditional path and the conditional source distribution. Instead, $\mathbf{h}_{i}$ is a condition of position $i$, also a time-invariant condition of the whole flow-matching process $t\in[0,1]$. Technically, we implement the flow-matching network by a small MLP:
@@
This procedure generates a predicted sample $\widehat{\mathbf{y}}_{i}$ at position $i$. To facilitate probabilistic forecasting, the procedure can be repeated using various initial noises, thereby enabling the computation of various statistics such as the median and quantiles from a set of generated probable predictions.
@@ 4.2 TimeBench
We collected and curated TimeBench, which comprises over 1 trillion time points from various sources, as shown in Figure 3. Several datasets originate from research teams (Woo et al., 2024; Ansari et al., 2024; Liu et al., 2024a, b). While most datasets are collected from real-world records, a small portion (0.05%) is generated synthetically to enhance pattern diversity, following KernelSynth proposed by Ansari et al. (2024). We also leverage substantial meteorological data (Hersbach et al., 2020) because of the predictability of weather systems. Data of different frequencies implies common and comprehensive temporal dynamics.
@@ 5 Experiments
We extensively evaluate Sundial on zero-shot forecasting benchmarks (Section 5.1) and investigate the scaling behavior across different model sizes (Section 5.2). We validate the effectiveness of TimeFlow compared to other training objectives (Section 5.3). We discuss the performance/speed trade-off during inference (Section 5.4) and conduct model adaptation to justify transferability of Sundial (Section 5.5).
Figure 4: Evaluation summary of Sundial.
@@
Table 2: Aggregated performance on GIFT-Eval, which comprises $23$ datasets characterized by a variety of frequencies, number of variates, and prediction lengths. We evaluate zero-shot forecasting performance using Sundial (Base). A lower MASE or CRPS indicates a better prediction. Rank assigns a numerical ranking of all $97$ configurations. Results of baselines are officially reported by Aksu et al. (2024).
@@
Model Naïve Seasonal Auto Auto DeepAR TiDE NBEATS PTST. iTrans. TimesFM VisionTS Chronos Moirai Sundial
Naïve ARIMA Theta (2020) (2023a) (2019) (2022) (2023) (2023b) (2024) (2024) (2024) (Ours)
MASE 1.260 1.000 0.964 0.978 1.206 0.980 0.842 0.762 0.802 0.967 0.775 0.786 0.809 0.727
CRPS 1.383 1.000 0.770 1.051 0.721 0.652 0.689 0.496 0.524 0.575 0.638 0.551 0.515 0.505
Rank 23.392 21.598 17.464 19.928 15.381 14.856 17.309 7.680 8.580 11.443 16.474 11.247 7.845 9.536
@@ 5.1 Time Series Forecasting
In this section, we focus on zero-shot forecasting, we compare Sundial with advanced time series foundation models on various benchmarks, including (1) point forecasting: we adopt the long-term forecasting benchmark (Wu et al., 2022), which assesses the performance under different forecasting horizons using MSE and MAE; (2) probabilistic forecasting: we experiment on GIFT-Eval (Aksu et al., 2024) and FEV leaderboard (Ansari et al., 2024), following their official evaluation suite and assessing point (MASE) and probabilistic (CRPS and WQL) metrics. All evaluated datasets are excluded from the pre-training dataset. The model configurations are detailed in Table 5.
@@ 5.1.1 Point Forecasting
As shown in Table 1, Sundial consistently outperforms other advanced time series foundation models. Compared with the previous state-of-the-art model Time-MoE (Shi et al., 2024b), the Sundial family using fewer parameters achieves the average MSE reduction of $7.57\%$ and averaged MAE reduction of $4.71\%$. Notably, continuous tokenization allows our model to conduct patch-level forecasting with fewer autoregression steps, while Chronos using point-wise discrete tokenization may not be effective in long-term forecasting.
@@ 5.1.2 Probabilistic Forecasting
Beyond point forecasting, Sundial possesses a unique generative capability for making probabilistic predictions. Following Ansari et al. (2024), we calculate the median and quantiles using $20$ generated raw predictions of Sundial. While several baseline models have been pre-trained by the consistent objective function for probabilistic evaluation, e.g., quantile loss for WQL, Sundial calculates these statistics for evaluation without any prior knowledge.
@@
We evaluate our Sundial on the open leaderboard established by AutoGluon (Ansari et al., 2024), which includes $27$ datasets for probabilistic forecasting. As shown in Figure 5, the zero-shot forecasting performance of Sundial exceeds $60\%$ statistical methods and deep models that are superwisedly trained in distribution. While Sundial is ranked as the second zero-shot pre-trained models after Chronos, Sundial realizes 11.34×11.34$\times$11.34 × inference time speedup as shown in Figure 6. Based on native patching and multi-patch prediction, our inference time is near to N-BEATS.
Qualitative showcases for point and probabilistic forecasting are presented in Appendix D. Our model generates highly eventful and coherent temporal patterns with input series.
@@ 5.2 Scalability
From Table 1, the larger Sundial model generally achieves better performance and more wins with the scaling of parameters. Beyond downstream performance, we delve into the utilization of model capacity. We plot training curves in Figure 7. Compared with Sundial (Small), the large version leads to totally $15.38\%$ reduction in the converged training objective, showing significant performance promotion on in-distribution time series forecasting.
@@ 5.3 TimeFlow Loss
Based on the flow-matching framework, TimeFlow Loss allows autoregressive models to learn and generate flexible distributions based on learned representations. To validate the effectiveness of this design, we implement two alternatives: (1) an MLP network and MSE Loss and (2) a parameterized training objective based on the denoising diffusion procedure (Li et al., 2024). We adopt the same parameterized network and Transformer backbone and pre-train them on TimeBench. Since the converged training loss is not comparable across different objective functions, we compare zero-shot performance in Table 3. Despite allowing for prediction generation, performance using diffusion-based objective is notably inferior to TimeFlow Loss.
@@
In addition to zero-shot performance, we provide showcases for quality evaluations in Appendix D.2. Pre-trained model optimized by the specific MSE Loss can only output a single prediction and the prediction is sometimes over-smooth due to mode collapse in large-scale pre-training. During pre-training, generative modeling can accommodate significantly different future variations even if their lookback series are similar. It benefits downstream tasks by generating multiple plausible predictions, which indicates various possibilities in time series forecasting, thereby facilitating the applicability for decision-making.
5.4 Model Inference
Generative modeling introduces flexibility to adjust predictions of pre-trained models during inference. Even with the naïve strategy to generate predictions, i.e., sampling different noise from a standard Gaussian distribution, there are two configurations to control the quality of predictions: (1) the number of sampled predictions to calculate statistics and (2) sampling steps specified in flow-matching. We present results with varied configurations in Figure 8.
@@ 5.4 Model Inference
The top two figures conform to the central limit theorem. Generating more samples consistently leads to more precise estimation of statistics and better results. The bottom two figures indicate that using fine-grained steps during the push-forward process generally leads to good predictions.
These observations reveal the trade-off between inference speed and performance, which provides flexibility for various use cases requiring different certainty in predictions. Note that using different configurations does not require retraining the model. In our experiments, we consistently adopt the choice of sampling $20$ predictions with each generated by $50$ steps, achieving comparable inference time as task-specific deep models as shown in Figure 6. Advanced strategies of sampling and post-processing of raw prediction leave interesting directions for future exploration.
@@ 6 Conclusion
In this work, we collect and curate TimeBench, a trillion-scale time series dataset for building time series foundation models, which can benefit the research community. Towards time series foundation models, we delve into tokenization and optimization, presenting contributions in two aspects. First, we demonstrate that continuous tokenization, such as patch series, can be more effective and efficient for the time series modality, and generative modeling presents a native approach for learning on continuous-valued time series. Second, we propose a novel training objective to accommodate heterogeneous time series distribution while endowing autoregressive models with an inherent capability to sample from non-categorical distribution. Our pre-trained Sundial models make substantial advances on best-recognized forecasting leaderboards. We hope this work can inspire future paradigms for pre-training time series foundation models and enhance their applicability to real-world scenarios.
@@ Appendix A Dataset Statistics
Large-scale datasets are of paramount importance for pre-training foundation models. Recent research has contributed significant time series datasets (Das et al., 2023b; Liu et al., 2024b; Shi et al., 2024b). While the scaling law of time series foundation models has been explored in the recent work (Shi et al., 2024a), the pre-training scale remains relatively limited. Given the heterogeneity of time series compared to other modalities, it raises the question of whether it is feasible to learn from enormous series. To address the question, we curated TimeBench including 1 trillion time points from various domains.
The statistical details of TimeBench are summarized in Table 4. In addition to open-source datasets from research teams on time series foundation models (Woo et al., 2024; Ansari et al., 2024; Liu et al., 2024b, a), we collected substantial real-world time series from various domains such as finance, IoT, meteorology (Muñoz-Sabater et al., 2021), and healthcare (Goldberger et al., 2000). These resources enable us to construct large-scale time-series corpora exceeding 1 trillion time points. The corpora include highly credible and predictable data with a wide range of frequencies, lengths, and number of variates, providing comprehensive temporal dynamics and variation patterns to facilitate downstream applications. To prevent data leakage, we exclude all datasets evaluated in Section 5.1 to make sure that Sundial conducts zero-shot forecasting.
Table 4: Key statistics of TimeBench, the pre-training dataset of Sundial, which encompasses various sources.
@@ Appendix B Implementation Details
All experiments are implemented using PyTorch (Paszke et al., 2019) and executed on NVIDIA A100 GPUs. We employ the AdamW optimizer (Kingma & Ba, 2014) for model optimization. We adopt channel independence (Nie et al., 2022) for univariate pre-training. During training, data from different domains is sampled according to a predefined ratio to balance the data volume across domains and ensure diversity in the training data. We implement a global shuffle strategy by loading time series into a standard parquet format. We use variable-wise normalization to unify the scope of values.
On the FEV leaderboard (Ansari et al., 2024), which consists of short-term forecasting datasets with a maximum prediction length of $56$, we train Sundial models by TimeFlow Loss with the prediction length of $F=16$. For the point forecasting (Wu et al., 2022) and GIFT-Eval (Aksu et al., 2024), which consist of forecasting datasets with a prediction length ranging from $6$ to $900$, we train Sundial models by TimeFlow Loss with the prediction length of $F=720$. For the required prediction length less than the model prediction length, we truncate the output generated by Sundial. For the required length more than the prediction horizon, we conduct rolling forecasting. Following the generative forecaster Chronos (Ansari et al., 2024), we sample $20$ raw predictions with each generated by $50$ sampling steps to calculate metrics for evaluation, including MASE, CRPS, and WQL. Detailed configurations of Sundial in different sizes are provided in Table 5. We provide a model summary in Table 6, which includes more counterparts and summarizes several essential aspects of time series foundation models.
@@
Method Sundial Time-MoE Timer Moirai MOMENT LLMTime Chronos Lag-Llama TimesFM
@@
Model Size 32M 113M 29M 14M 40M - 46M 200M 17M
128M 453M 50M 91M 125M 200M 70M
 [^4] M 2.4B 67M 311M 385M 710M 200M
Pre-training Scale 1034B 300B 231B 231B 1.13B - 84B 0.36B 100B

### v2

@@ 4.1.3 TimeFlow Loss
Given representations $\{\mathbf{h}_{i}\}$ extracted by the last layer of the Transformer, we aim to generate length-$F$ predictions $\widehat{\mathbf{y}}_{i}=\widehat{x}_{1+iP,F+iP}$ at each position $i$ via our autoregressive model. Motivated by the empirical observation that a larger patch size improves the performance in decoder-only Transformers (Das et al., 2023b) while a small patch size can be more flexible to accommodate data of different frequencies, we adopt multi-patch predictions ($F>P$) for pre-training, which also reduces the steps of autoregressive inference.
@@
where $\mathbf{y}_{i}\in\mathbb{R}^{F}$ is the groundtruth value and $\mathbf{y}_{i}^{(0)}$ is a $d$-dimensional Gaussian noise, $t$ is sampled from $\mathcal{U}[0,1]$, and $\mathbf{y}_{i}^{(t)}=t\mathbf{y}_{i}+(1-t)\mathbf{y}_{i}^{(0)}$ is constructed by the conditional optimal-transport path. It is important to note that the conditional representation $\mathbf{h}_{i}$ differs from the conditional path and the conditional source distribution. Instead, $\mathbf{h}_{i}$ is a condition of position $i$, also a time-invariant condition of the whole flow-matching process $t\in[0,1]$. Technically, we implement the flow-matching network by a small MLP:
@@
This procedure generates a predicted sample $\widehat{\mathbf{y}}_{i}$ at position $i$. To calibrate probabilistic forecasting results during inference, we repeat this procedure using different initial noises and estimate statistics such as the median and quantiles from a set of generated predictions. We implement an efficient repeated-sampling in the TimeFlow module. The condition (representation) of lookback series is shared and reused for different initial noises, thereby reducing the overhead of repeated forwarding in the Transformer backbone.
@@ 4.2 TimeBench
We collected and curated TimeBench, which comprises over a trillion time points from various sources, as shown in Figure 3. Several datasets originate from research teams (Woo et al., 2024; Ansari et al., 2024; Liu et al., 2024a, b). While most datasets are collected from real-world records, a small portion (0.05%) is generated synthetically to enhance pattern diversity, following KernelSynth proposed by Ansari et al. (2024). We also leverage substantial meteorological data (Hersbach et al., 2020) because of the predictability of weather systems. Data of different frequencies encompasses common and comprehensive temporal dynamics.
@@ 5 Experiments
We evaluate Sundial on best-recognized zero-shot forecasting benchmarks (Section 5.1) and investigate the scaling behavior of Sundial (Section 5.2). We compare TimeFlow with other training objectives (Section 5.3). We delve into test-time calibration of generative forecasters (Section 5.4). We conduct model adaptation of Sundial, i.e., instruction tuning (Section 5.5) and provide in-depth ablation studies to evaluate our modular enhancement (Section 5.6).
@@
Table 2: GIFT-Eval comprises $23$ datasets characterized by a variety of frequencies, variate numbers, and prediction lengths. We evaluate zero-shot performance using $100$ generated series, being consistent with Woo et al. (2024). A lower MASE or CRPS indicates a better performance. Rank assigns a numerical ranking of all $97$ configurations. Baseline results are officially reported by Aksu et al. (2024).
@@
Model Naïve Seasonal Auto Auto DeepAR TiDE N-BEATS PTST. iTrans. TimesFM TabPFN Chronos Moirai Sundial
Naïve ARIMA Theta (2020) (2023a) (2019) (2022) (2023a) (2023b) (2025) (2024) (2024) (Ours)
MASE 1.260 1.000 0.964 0.978 1.206 0.980 0.842 0.762 0.802 0.680 0.748 0.786 0.809 0.673
CRPS 1.383 1.000 0.770 1.051 0.721 0.652 0.689 0.496 0.524 0.465 0.480 0.551 0.515 0.472
Rank 28.072 26.175 21.515 24.031 18.938 18.557 21.381 10.052 11.320 8.237 8.268 14.309 10.175 9.062
@@ 5.1 Time Series Forecasting
In this section, we focus on zero-shot forecasting, we compare Sundial with advanced time series foundation models on various benchmarks, including (1) point forecasting: we adopt the long-term forecasting benchmark (Wu et al., 2022), which assesses the performance under different forecasting horizons using MSE and MAE; (2) probabilistic forecasting: we experiment on GIFT-Eval (Aksu et al., 2024) and FEV leaderboard (Ansari et al., 2024), following their official evaluation suite and assessing point (MASE) and probabilistic (CRPS and WQL) metrics. All evaluated datasets are excluded from the pre-training dataset. Model is available on HuggingFace [^1] https://huggingface.co/thuml/sundial-base-128m. and configurations are detailed in Table 5.
@@ 5.1.1 Point Forecasting
As shown in Table 1, Sundial consistently outperforms other advanced time series foundation models. Compared with the previous state-of-the-art model Time-MoE (Shi et al., 2024b), the Sundial family using fewer parameters achieves the average MSE reduction of $7.57\%$ and averaged MAE reduction of $4.71\%$. Notably, continuous tokenization allows our model to conduct patch-level forecasting with fewer autoregression steps, while Chronos using point-wise discrete tokenization may not be suitable in long-term forecasting.
@@ 5.1.2 Probabilistic Forecasting
Beyond point forecasting, Sundial possesses a unique generative capability for making probabilistic predictions. Following Ansari et al. (2024), we calculate the median and quantiles using a set of raw predictions of Sundial. While several baseline models have been pre-trained by the consistent objective function for probabilistic evaluation, e.g., quantile loss for WQL, Sundial calculates these statistics for evaluation without any prior knowledge.
@@
We evaluate our Sundial on the open leaderboard established by AutoGluon (Ansari et al., 2024), which includes $27$ datasets for probabilistic forecasting. As shown in Figure 4, the zero-shot forecasting performance of Sundial exceeds $70\%$ statistical methods and deep models that are superwisedly trained in distribution. While Sundial is ranked as the second zero-shot pre-trained models after Chronos, Sundial realizes 35×35$\times$35 × inference speedup as shown in Figure 5. Based on patch-wise tokenization and multi-patch prediction, our inference speed is near to N-BEATS.
Besides, we provide qualitative showcases in Appendix D. TimeFlow can generate highly eventful and coherent temporal patterns with input series. Beyond the mean or quantiles, our model enables the estimation of arbitrary statistics by sampling directly from the predictive distribution.
@@ 5.2 Scalability
From Table 1, the larger Sundial model consistently achieves better performance with the scaling of parameters. Beyond downstream performance, we delve into the utilization of model capacity. Figure 6 shows training curves of different sizes. Compared to Sundial (Small), the large version leads to $15.38\%$ reduction in the converged training loss, exhibiting promising model capacity of generative forecasters.
@@ 5.3 TimeFlow Loss
Based on the flow-matching framework, TimeFlow Loss allows autoregressive models to learn and generate flexible distributions while enhancing representation learning. To validate the effectiveness of this design, we implement two alternatives: (1) an MLP network and MSE Loss and (2) a parameterized training objective based on the denoising diffusion procedure (Li et al., 2024). We adopt the same parameterized network and Transformer backbone and pre-train them on TimeBench. Since the converged training loss is not comparable across different objective functions, we compare zero-shot performance in Table 3. Despite allowing for sampling predictions, performance using diffusion-based objective is notably inferior to TimeFlow Loss.
@@
In addition to zero-shot performance, we provide showcases for quality evaluations in Appendix D.2. Pre-trained models optimized by the specific MSE Loss can only output a single prediction. And the prediction is sometimes over-smooth due to mode collapse (refer to Appendix C.1). Instead, generative modeling can accommodate significantly different future variations even if their lookback series are similar. We also provide a probablistic metric CRPS to compare different objectives in Table 7, which validate that the predictive distribution modeled by TimeFlow is more coherent and diverse than counterpart training objectives. It benefits downstream tasks by generating multiple plausible predictions, conveys various future possibilities and enhances the reliability of decision-making.
5.4 Test-Time Calibration
Generative modeling facilitates the flexibility to calibrate the final prediction during inference. Based on the median-based forecasting strategy, i.e., starting from multiple noise of a standard Gaussian and calculating the median of raw predictions, there are two configurations to calibrate final predictions: (1) the number of samples to calculate statistics and (2) sampling steps $K$ used for flow-matching. Figure 7 shows the results using different configurations.
@@ 5.4 Test-Time Calibration
The top two figures conform to the central limit theorem. Generating more samples leads to more calibrated estimation of prediction and confidence interval. The bottom two figures indicate that using fine-grained steps during the push-forward process can leads to more precise predictions.
The trade-off between inference time and performance reveals the potential of test-time calibration, which does not require retraining models. The generative capability of Sundial provides flexibility for various use cases requiring different levels of uncertainty. In our experiments, sampling $20$ predictions with each generated by $50$ steps consumes nearly one second on a CPU, which is notably more efficient than tuning deep models or statistical methods. Advanced strategies of sampling and post-processing of raw prediction leave interesting directions for future exploration.
@@ 5.5 Model Adaptation
Figure 9: Ablation studies with respect to architectural enhancements. We report the averaged results of TSLib datasets (Wu et al., 2022) from four prediction lengths $\{96,192,336,720\}$ and all six datasets. The context length is set to $2880$ and the patch length is $16$.
5.6 Ablation Study
We conducted several ablation studies that provide insights into the enhancement made to Sundial’s architecture. We evaluate the overall zero-shot performance on TSLib, which covers six different datasets and four prediction lengths.
RoPE
Prior research (Liu et al., 2024a) observed that the introduction of RoPE (Su et al., 2024) yields better results in supervised forecasting tasks. As shown in Figure 9 (a), RoPE can also improve zero-shot forecasting, presenting a general enhancement for time series foundation models.
Layer Normalization
Pre-LN (Baevski & Auli, 2018) is widely adopted in large language models (Touvron et al., 2023) due to the training stability. As depicted in Figure 9 (b), training with Pre-LN for more iterations yield better performance. In contrast, training with Post-LN, which is the predominant choice in supervised models, may adversely affect downstream results.
FlashAttention and KV Cache
We make it to leverage FlashAttention (Dao et al., 2022) and KV Cache to reduce the computational costs. As shown in Figure 9 (c) and (d), they notably reduce $14.8\%$ memory footprint and $43.6\%$ inference time without affecting performance.
@@ 6 Conclusion
In this work, we collect and curate TimeBench, a trillion-scale time series dataset for building time series foundation models, which can benefit the research community. Towards time series foundation models, we delve into tokenization and optimization, presenting contributions in two aspects. First, we demonstrate that continuous tokenization, such as patch tokens, can be more effective and efficient for the time series modality, and generative modeling presents a native approach for learning on continuous-valued time series. Second, we propose a novel training objective to accommodate heterogeneous time series distribution. It endows autoregressive models with an inherent capability to sample from non-categorical distribution. Our pre-trained Sundial models make substantial advances on best-recognized forecasting leaderboards. We hope this work can inspire future paradigms for pre-training time series foundation models and enhance their applicability to real-world applications.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (U2342217 and 62021002), the BNRist Innovation Fund (BNR2024RC01010), and the National Engineering Research Center for Big Data Software.
We extend our gratitude to Xingzhuo Guo for his expertise in flow-matching and meticulous proofreading of the method section. We further thank Jialong Wu, Yuezhou Ma and Yu Zhang for insightful discussions about generative models. Their collective support significantly enhanced this work.
Impact Statement
This paper aims to advance the development of time series foundation models. We curated a pre-training dataset from publicly available resources. Our models employ an efficient tokenization and incorporate a generative training objective. The proposed TimeFlow Loss provides insights for training generative foundation models for time series forecasting. We released our pre-trained models that demonstrate notable zero-shot forecasting performance and versatility across different benchmarks. The generative forecasting paradigm enhances the model reliability for decision-making. The millisecond inference speed of our model can be beneficial to many practitioners. Our paper mainly focuses on scientific research and has no obvious negative social impact.
@@ Appendix A Dataset Statistics
Large-scale datasets are of paramount importance for pre-training foundation models. Recent research has contributed significant time series datasets (Das et al., 2023b; Liu et al., 2024b; Shi et al., 2024b). While the scaling law of time series foundation models has been explored in the recent work (Shi et al., 2024a), the pre-training scale remains relatively limited. Given the heterogeneity of time series compared to other modalities, it raises the question of whether it is feasible to learn from enormous series. To address the question, we curated TimeBench with a trillion time points from various domains.
Unlike other modalities, most time series are unavailable on open websites or repositories. There are also limited domains that encompass typical and predictable time series, leading to slow progress on dataset construction. Therefore, we conducted tedious preprocessing, including missing values imputation, abnormalities exclusion, and normalization techniques. We conducted statistical analysis, examining time series through the lenses of intrinsic properties, e.g., non-stationarity, forecastability, and seasonality. This approach allows us to characterize the data quality inherent to time series, which affects the training stability of next-token prediction. We also adopt synthetic techniques to improve pattern diversity. Further, we adopt ERA5 (Muñoz-Sabater et al., 2021), including systematic real-world temporal observations.
The statistical details of TimeBench are summarized in Table 4. In addition to open-source datasets from research teams on time series foundation models (Woo et al., 2024; Ansari et al., 2024; Liu et al., 2024b, a), we collected substantial real-world time series from various domains such as finance, IoT, meteorology, and healthcare (Goldberger et al., 2000). These resources enable us to construct large-scale time-series corpora exceeding a trillion time points. The corpora include highly credible and predictable data with a wide range of frequencies, lengths, and numbers of variates, providing comprehensive temporal dynamics and variation patterns to facilitate downstream applications. To prevent data leakage, we exclude all datasets evaluated in Section 5.1 to make sure that Sundial conducts zero-shot forecasting.
Table 4: Key statistics of TimeBench, the pre-training dataset of Sundial.
@@ Appendix B Implementation Details
All experiments are implemented using PyTorch (Paszke et al., 2019) and executed with $32$ NVIDIA A100 GPUs. We employ the AdamW optimizer (Kingma & Ba, 2014) for model optimization. We adopt S3 format (Liu et al., 2024b) for univariate pre-training. During training, data from different domains is sampled according to a predefined ratio to balance the domain weightings and ensure diversity in the training data. We implement a global shuffle strategy by loading time series into a standard parquet format. We use variable-wise normalization to unify the scope of values.
On the FEV leaderboard (Ansari et al., 2024), which consists of short-term forecasting datasets, we train Sundial models by TimeFlow Loss with the prediction length of $F=16$. For the point forecasting (Wu et al., 2022) and GIFT-Eval (Aksu et al., 2024), which consist of forecasting datasets with a prediction length ranging from $6$ to $900$, we train Sundial models by TimeFlow Loss with the prediction length of $F=720$. For the required prediction length less than the model prediction length, we truncate the output generated by Sundial. For the required length more than the prediction horizon, we conduct rolling forecasting. Following Chronos (Ansari et al., 2024), we sample $20$ raw predictions to calculate MASE and WQL on FEV. Being consistent to Moirai (Woo et al., 2024), we sample $100$ raw predictions to calculate MASE and CRPS for GIFT-Eval. The sampling step is fixed as $K=50$. Configurations of Sundial in different sizes are provided in Table 5. We provide a model summary in Table 6, which summarizes several aspects of current time series foundation models.
@@
Method Sundial Time-MoE Timer-XL Moirai MOMENT LLMTime Chronos Lag-Llama TimesFM
@@
Model Size 32M 113M 84M 14M 40M - 46M 200M 17M
128M 453M 91M 125M 200M 70M
 [^4] M 2.4B 311M 385M 710M 200M
Pre-training Scale 1032B 300B 260B 231B 1.13B - 84B 0.36B 100B